OpenStudy (anonymous):

A block of mass 30 kg is being brought down by a chain. If the block acquires a speed of 40cm/s in dropping down 2m. The work done by the chain during the process will be..??

OpenStudy (anonymous):

Is answer 24joules

OpenStudy (anonymous):

\[E = 1/2\times Mass \times Velocity^2\] Work done is the energy transferred during an action, so the work done during this example is equal to the kinetic energy transferred to the brick by the time it's finished moving those 2m. So, working out the kinetic energy from the equation above (E= 1.2 x ...) will give you the work done. Inputting the data you provided: \[E = 1/2 \times 30 \times (0.40)^2\] I've left the units out of the equation and have converted the speed of 40cm/s into metres/second to give 0.40 m/s. Working out this equation gives= 2.4J Thus, the kinetic energy of the brick, and thus the work done, to the brick whilst it has moved the 2m is 2.4J. Hope this helps!